∫¹∫2 (cotx + sinx)dx


1
Expert's answer
2020-04-30T19:49:07-0400

"\int_{1}^{2}(cotx+sinx)dx=(\ln sinx-cosx)""|_1^2=\ln sin2-\ln sin1-cos2+cos1\approx1.03"